Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

Actions

  • [van Moorsel]: Outline OSRO observation checklist (all required steps in delivery of archive data to PI).
  • [McMullin]: Add ECSO Projects to JIRA/tracking.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Description of guideline reduction steps in continuum processing.
  • [Perley/Dhawan]: Develop test to develop/demonstrate readiness for high frequency observing.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Provide an update to the default frequencies for L band. Note: These are for 128 MHz (64 MHz could be different). FO asked if there is a lower frequency setting that could be used on the Low/High mode (shift above L3 GPS at 1381). Procedure currently is to notify the PI for approval. Rick noted that he has 250 kHz resolution spectra from 1 - 2 GHz; put this in the public domain and let it be reviewed by the proposers.
    • Low/High: 1452 MHz, 1860 MHz
    • Two low: 1328 MHz, 1456 MHz
    • Two high: 1732 MHz, 1860 MHz
  • [Perley]: Perform a frequency sweep test at S band to provide to engineers for review (Q2).
  • [Benson]: Update when automated handling of multiple spectral windows is available (Q2). Clarified: John will organize a small group this week (likely necessary candidates are Eric, Michael, Rick) to resolve the near and long term plans for handling the automated export to UVFITS for heterogeneous spectral windows (e.g., auto split by spectral window, etc). Results to be presented next Monday.

Commissioning review/reports

News

  • Leap forward in system scheduling; able to chain together SBs without requiring the CBE up/down intervention. Many projects executed over the weekend.
  • Note that there were problems with the data from the weekend (many zero amplitude visibilities). Please review your data over the last week for this and summarize the fraction of data lost in this way (previously this was less than 1% of the data).
  • Bob, Steve (Myers), Steve (Durand), Jeff (Kern) (+ others from CV) at International SKA Science and Engineering Meeting
  • EVLA Test Meeting (Rick); will review recent results and evaluate those that can/require an in depth discussion; these will be scheduled for the monthly Test Meeting. Rick will propose topics and ask for volunteers for areas where detailed/demonstrable results can/should be discussed with the engineering/science/etc gurus. Rick/Joe will also review the lists of JIRA tickets for possible topics; this once monthly test will not be able to cover all details of testing but will focus on in-depth discussions of a subset as needed. Rick will chair this meeting.

Notes for week of 22 Mar 2010 (Black: old but still relevant; Blue: new)

  • Due to issues of conflict/slow down for multiple access to the same file, please wait for John to copy to the data holding area at: /home/evla-data and grab it from there.
  • Quick Software Systems review:
    • OPT (Lorant): Project codes are getting into the archive now (much easier to read; note: any projects that are edited will still have the lovely TOSR000n_sbxxxxxxx... in there). Update planned for Wednesday that will allow editing SBs directly (for super-users). Resolution to X band tuning issues.
    • Executor (Barry): Introduced a bug in the reference pointing; expected to be resolved in the next day or two.
    • Scheduling (Joan, Bryan): Test of OST today at 1500.
    • CASA (Juergen): importevla available in casapy-test. Michael noted the issue of the performance of time averaging in split (brought up at the OSRO commissioning testers meeting). Information on timings for data sets of different sizes and observing times should be noted and sent to Gustaaf and Juergen (Juergen will note this at the weekly CASA meeting). Still tracking (and need subsequent information) on the "large" data set import failures (seen with AE176). Send any instances to Gustaaf/John.
    • AIPS (Eric): No AIPS meeting this Wednesday.
  • Hardware notes (Steve, Ken)
    • Delay errors - largest at the 10 ns level; pending return of EA24 (used for single dish/receiver testing), update will be scheduled. Ken has taken the data already.
    • EA16 X band RCP receiver is dead (no power to down converter).
    • EA05/EA01 follow-up; observations from AE176 indicate that EA05 is not the only problem at L band; EA01 also has issues.
    • EA07 is weak at all 3 bands. C, X, K.
    • EA22 needs pointing measured at K, Ka, Q.
    • EA03 A/C IF is dead (works only at default C band); to be replaced today. Replaced - please comment if there are any issues.
    • EA17 has passed electronics acceptance tests (K, X, C are functioning on all four IFs). Included in all observations over the weekend.

OSRO Priority Testing Status

  • OSRO Readiness Master List
        • Updated weekly based on the Monday morning meeting. Note that lower priority LCX are now getting scheduled due to the delays in enabling high frequency observing. Key Priority 0 items covered below.

System Verification

  • WIDAR update (Michael)
    • Things are more automated now (can prepare a project in the OPT, write it out and pass it to the operator for observing at a given time). Projects can have a single correlator configuration per SB; multiple resource settings currently require modifications to the script and so are not being used. Goals are to: 1) handle more than 1 correlator configuration (hopefully tomorrow for testing), 2) utilize the scheduler, 3) push to modes beyond OSRO (currently only using 4 BB; more are available; similarly, we are only using one CBE notde (currently 3 are available) - we should be able to get up to 1 GHz BW. This week the first version of the final switch (which moves data between the BB and the CBE) will be tested.
    • Ken/Adam are working on the implementation/testing of the phased array; note: Adam leaves Socorro mid-May.
    • Ken is working on getting/testing switched power out of the BB.
    • Status of dummy scans: previously needed two dummy scans (1 minute): 1) set attenuator levels, 2) bugs in the CBE required an additional scan to avoid throwing away data; the need for this has gone away (so only need one dummy scan now). It was noted that the executor gets the phase wrong on the 1st scan but this can be the dummy scan when the attenuators are being set. Frazer asked about fast switching - will this require 2 dummy scans? Michael: Yes, one per frequency setup.
    • Issues with subbands less than 500 kHz (which use the last stage of the filters) have issues which have been reproduced; waiting on fix.
  • Receiver update (Rick): Spent several days at the site; most has been summarized in notes to evlatests but will write up more.
    • Note effort to summarize tuning capabilities in: https://bugs.aoc.nrao.edu/browse/EVL-1137
    • S/C band RFI tests (Rick)
      • Bob Hayward and I did some tests a year ago on the prototype S-band system, using the 'Bandpass Tool', which provides good frequency resolution over a 1 GHz-wide bandpass. I have reviewed these plots taken with a center frequency of 3800 MHz, (hence showing the spectrum from 3300 to 4300 MHz). The results are interesting ... The data taken with the antenna vertical, and pointing at Cyg A (high elevation) appear fine. But the 'tip' data, taken mostly at low elevation, show moderate level RFI, extending more or less uniformly between 3600 and 4200 MHz.
    • S Band Testing (Rick): Noted that S is quite bad for RFI; in particular satellite radio (2312-2348 MHz); also noted satellite downlinks at 2185. FSS (Fixed Satellite Services) from 3600-4200 used for downloading programs to affiliates. S band sensitivities and efficiencies are meeting expectations (20-30s and efficiencies of around 50 percent). Also noted that the spillover on S band at low elevation is high (S band horn is on the upper side of the feed so you can see the ground when pointed at the horizon).
    • Ku Band Testing (Rick): Looks great (Tsys in center is in the 20s, Trx ~ 4-5 K); continuing to reduce the data from the hot-cold/pointing rasters and sky dips.
    • L band, HI checkout (ECDATA-47); Juergen showed the results of using three tunings to enable adequate bandpass calibration. Will write up recommendation to be passed along.
  • High Frequency Readiness
    • Reference pointing status (Manuel/Vivek): Waiting for real test rather than C/K stop gap.
    • set-and-remember status (Rick): Issues during single antenna tests (excited a mode not anticipated - to be fixed). No issues.
      • within band frequency changes (ECDATA-33): okay (with antenna/subband exceptions).
        • Will see change until LO fiber is included in antenna phase model (Rick?)
  • Gain transfer: ok
  • OSRO1 BW/channelization Tests (ECDATA-70; Vivek)
    • Overall worked great. No fine scale structure (so can use wide band to calibrate narrow band). Polarization as expected (same D for different bandwidths). Michael noted the 4 stage digital filters (1st for 128 MHz, 2nd for 8-64 MHz, 3rd for 0.5 - 4 MHz); see some digital filter effects at 4 MHz bandwidths; these can be taken out. Frazer noted that we want to analyze the stability and so should re-run a week later to see.
  • OSRO2 BW/channelization Tests (ECDATA-75; Vivek)
    • Will be looked at.

Band Checkouts

  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated. Okay?
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40): Update from Frazer.
        • James reviewed the data further - found that different channel averagings did integrate down; others not so much; still an issue to be explored overall. Frazer is looking at the field and will present results soon.
  • S Band Status: Ready. S band edges test data was taken to see if any effects like at C bands low edge; to be reviewed.
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

Mode testing

  • Polarimetry Status:
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Eric reviewing AH1004 data for AIPS
    • George will be pushing things into JIRA shortly; noted that he is seeing an East/West difference in the data, don't settle to the same value; good candidate for a more detailed discussion (e.g., ionospheric effects, etc).
    • *Rick noted that the design for L, S, and C cross bar will be much better balanced and should show lower cross polarization.
  • Spectral Line

OSRO Proposal Testing. Gustaaf noted that they have a prototype checklist for review of projects that is in use; still being refined/evolving.

  • AE176 2nd run ([[https://bugs.aoc.nrao.edu/browse/ECDATA-71][ECDATA-71]): Gustaaf, Jacqueline
  • AK730 triggered GRB observations ([[https://bugs.aoc.nrao.edu/browse/ECDATA-72][ECDATA-72]): Joe, Dale
  • AY 199 HI ([[https://bugs.aoc.nrao.edu/browse/ECDATA-73][ECDATA-73]): Emmanuel
  • AS1012 ([[https://bugs.aoc.nrao.edu/browse/ECDATA-74][ECDATA-74]): Juergen
  • AU129 ([[https://bugs.aoc.nrao.edu/browse/ECDATA-77][ECDATA-77]): Gustaaf
  • AR704 - rerun
  • DEMO SCIENCE ([[https://bugs.aoc.nrao.edu/browse/ECDATA-xx][ECDATA-xx]): James
    • 1st 4 hours are okay but then the operators began taking/replacing antennas in the array (Joan: How do we not do this again; Michael: We ask them not to).

    • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • High Priority [Pokorny]: Auto-controlled CBE writes scrambled visibilities.
      • High Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [Butler]: Scan labeling synchronization.
      • Med Priority [Perley]: Issues integrating down when averaging by channel.
      • Med Priority [Myers]: C band 128 MHz spectrum bends down at the edges.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). EA03 R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Also seen on EA07, EA25 (both RR and LL). EA03 should be fixed.
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

Commissioning plans for the week

System Tests

    • [Rick]: Enable High Frequency Observing: principally the reference pointing confirmation for C/K; understanding of the set-and-remember good/bad antennas to employ for calibration. If weather permits, test K band this week.
    • [Rick]: Enable RSRO Mode Observing
    • [Rupen; OSRO Priority 0]: OPT-> Archive path complete (currently just automated setup of the correlator).End of week possible
    • OSRO-1 (256 MHz): Still need tests of polarization phase convention
    • OSRO-2: Review fringes, gain/bandpass stability, phase sign, frequency labeling at L band.
    • OSRO-1,2: Review behavior under bandwidth options; change BW in OSRO 1 in steps of factor of 2 from 256 MHz -> 1 MHz and in OSRO 2 in steps of factor of 2 from 256 MHz -> 250 kHz

OSRO project tests

As possible; see overview at: https://staff.nrao.edu/wiki/bin/view/EVLA/OSROReadiness_Cycle1; Note: Once automated corr setup is enabled; open up the ECSV plan observations (see planning review).
  • Tool from Lorant to see available SBs: /home/mchost/evla/bin/sbStatReport
    • AU129
    • AM1009
    • AR704
    • AV317
    • More projects to be updated throughout the week.

RSRO

  • Observations
  • Residents
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Adam Deller (POC: Bryan Butler; area(s): Correlator checkout)
    • James Miller-Jones (POC: Claire Chandler; area(s): Demo Science)
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests)
    • Dave Green (arriving; POC: Sanjay Bhatnagar); area(s): Imaging

This week in JIRA

  • What's the difference between resolved and closed?
    • The intention is that the person assigned to the action/task is able to resolve it (based on the description of what is required). It is intended to be closed by the submitter to indicate that the resolving solution satisfies the full intent of the submitted action.
  • What is expected in the resolution?
    • Key questions addressed (quantitatively). Post any images/spectra; post script of reduction as possible.

AOB

Planning review (for Q1/Q2 2010):

Followup deferred. Begin to look at the detailed needs for the tests being scoped out by the leads. Want to finalize the detailed observing/testing plans by the end of this month and have a view of the longer term commitments needed. Establish an overview of the obstacles to deploying to the community; in particular, understand all of the details of the end-to-end data flow/processing to make decisions on what is working versus what is usable by the community. Something like:

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
Example limit of 1 s int ok OPT selected limits not possible no recommendations avail ok can't cook coffee weather table - transient issue no-go

  • Critical Commissioning
    • System Verification (Perley/Rupen): Need plan to conclude complex gain, bandpass stability tests at L-Q, reference pointing, gain curves, tipping scans.
    • Observing Band Checkout (principally calibration strategies, flagging)
    • Observing Mode Support (broad modes based on science goals)
    • Observing support (e2e capability)
      • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
      • Flux Calibrator Models (Amy); ECSV-43
  • Advanced Science Support

    • Need:
      • [Butler] Planning: Planetary observing (OPT setup; correlator needs); ECSV-7, ECSV-15
      • [Chandler] Planning: Demo Science
      • [Frail] Planning: ToO scheduling heuristics
      • [Sjouwerman] Planning: OPT tests for primary frequencies; ECSV-6
      • [van Moorsel] Planning: Data distribution plan
      • [Sjouwerman] Planning: Archive functionality tests; ECSV-11
      • [Mioduszewski] Planning: Calibrator survey; ECSV-27
      • [Brisken] Planning: Phased array (VLBI, pulsars); ECSV-22, ECSV-23
      • [Golap] Planning CASA multi-thread/cluster testing.
      • [Walker] Planning VLBI setup
-- JosephMcMullin - 21 Mar 2010
Topic revision: r7 - 2010-03-23, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback